OpenStudy (anonymous):

Given the rational function f(x) = (3x^2 -7x+2)/(x^2 +4x +3) find all asympototes and intercepts

OpenStudy (anonymous):

What is the limit of f(x) when x goes to infinity?

OpenStudy (anonymous):

\[ \lim_{x\to \infty} (3x^2 -7x+2)/(x^2 +4x +3)=3 \] Hence y=3 is horizontal asymptote.

OpenStudy (anonymous):

The denominator can be factored \[ x^2+4 x+3=(x+1) (x+3) \] Hence x=-1 and x=-3 are vertical Asymptotes.

OpenStudy (anonymous):

The numerator can be factored as \[ 3 x^2-7 x+2=(x-2) (3 x-1) \] hence f(x) intercepts the x-axis on x=2 and x=1/3

OpenStudy (anonymous):

If x=0, then y=2/3. that is the y-intercept.